Greenhab Report – March 26th

Crew 207 GreenHab Report – 26-Mar-2019

GreenHab Officer: Julielynn Wong

Environmental control: Ambient with door open

Average temperatures: 30.8 degrees Celsius

Low: 25 degrees Celsius

High: 36.6 degrees Celsius

Hours of supplemental light: None

Daily water usage for crops: 4 gallons

Water in Blue Tank—132 gallons

Time(s) of watering for crops: 1230; 1630

Changes to crops: None

Narrative: GHO used a large funnel from the Science Dome to prevent low back strain when filling the Greenhab watering can.

Harvest: Red leaf lettuce 4 grams

Support/supplies needed: None

Sol Summary – March 26th

Crew 207 Sol Summary Report 26-Mar-2019

Sol: 10

Summary Title: A Day of Rest and Solitude

Author’s name: Julielynn Wong

Mission Status: Nominal

Sol Activity Summary:

Crew 207 continued 3D printing in the Hab today. We began printing on our third 3D printer today.

The following items were identified as useful to 3D print at MDRS:

Funnels (small and large)

COM has taken over GHO duties. XO has taken over ENG duties.

Look Ahead Plan: Continue 3D printing in Hab. Complete data collection for bike study. Prepare for arrival of Crew 208.

Anomalies in work: None

Weather: Status

Crew Physical Status: Nominal

EVA: n/a

Reports to be filed: Operations, Greenhab

Support Requested: None

Astronomy Report – Mar 25th

Crew 207 Astronomy Report 25 Mar 2019

Crew Astronomer: Tom Baldwin


Robotic Telescope Requested: None

Objects Viewed:

Remote observatory used to image M 65. Image processing not yet complete.

Problems Encountered: None


Solar Features Observed: None.

Observatory closed, and photos of telescope, Laptop, and Solar Astronomy box sent to Peter.

Problems Encountered: None.

GreenHab Report – March 25th

Crew 207 GreenHab Report – 25 Mar 2019

GreenHab Officer: Diane Rothberg, Julielynn Wong

Environmental control: Heat and cooling w/ambient air 10:00 hr

Shade Cloth (40%)

Average temperatures: RH

Low: 17 deg C 16%

High: 27 deg C 28%

Hours of supplemental light: 0 Hours

Daily water usage for crops: 6 gal

Water in Blue Tank: 136 gallons

Crops watered at 08:00 hours (3 gal) and 17:00 hours (3 gal)


1) 8:00 AM 26 deg C / 25%

2) 10:35 AM 17 deg C / 20%

3) 14:10 PM 27 deg C / 20%

4) 17:00 PM 24 deg C / 20%

Changes to crops:

Narrative: Disposed of a dead pea plant from CRew 206.

Harvest: Basil: 6 gm Parsley 6 gm Sage 1 gm

Support/ Supplies needed: none

Operations Report – Mar 25th

Crew 207 Operations Report 25 Mar 2019

SOL: 9

Name of person filing report: Tom Baldwin

Non-nominal systems:

Generator: Radiator hoses and alternator need replacement on the generator. A contractor has been arranged to make these repairs within a few days.


Generator (hours run): 12.5 hours

Generator turned on (1930 last PM), SOC (100%)

Generator was turned off at (0700 this AM), SOC (100%)

Diesel Reading – 50 %
Propane Reading – 60 %
Ethanol Free Gasoline – 2 gallons.

Water (auxiliary tank) – Not in use
Water (static tank) – 200 gallons (replenished by David)
Auxiliary to Static tank transfer– Not applicable

Gallons transferred: Not Applicable

Water in Green Hab – 142 gallons
Water (loft) – Static to Loft Pump used – yes
Water Meter: 142050.1 gallons

Toilet tank emptied: No

Deimos rover used: No, not functional

Hours: N/A

Beginning charge: N/A

Ending charge: N/A

Currently charging: N/A

Sojourner rover used: ASSIGNED TO DIRECTOR

Hours: Not applicable

Beginning charge: Not applicable

Ending charge: Not applicable

Currently charging: Not applicable

Spirit rover used: Not used

Hours: 85.1 hours

Beginning charge: 100 %

Ending charge: 100 %

Currently charging: Yes

Opportunity rover used: Not used


Beginning charge: 100 %

Ending charge: 95 %

Currently charging: Yes (now 100%)

Curiosity rover used: Used

Hours: 84.7 hours

Beginning charge: 100 %

Ending charge: 97 %

Currently charging: Yes (now 100%)

Notes on rovers: not used

ATV’s Used: (Honda, 300, 350.1, 350.2, 350.3): N/A

Reason for use: Not used

Oil Added?: N/A
ATV Fuel Used: N/A

# Hours the ATVs were Used today: None

Notes on ATVs: None

HabCar : None.

CrewCar used and why, where? Not used

General notes and comments: Nothing to report

Summary of internet: Nothing to report

Summary of suits and radios: EVA suits normal, with all batteries greater than 13 volts,

Summary of Hab operations: Per recent electrical anomalies, checked SOC of electrical system.

06:00 SOC 100 %, generator on.

07:30 SOC 100%, generator off

10.50 SOC 91%

Summary of Green Hab operations: Nothing to report

Summary of ScienceDome operations: Nothing to report

Summary of RAM operations: Nothing to report

Summary of any observatory issues: Nothing to report

Questions, concerns and requests to Mission Support: None

Sol Summary Report – Mar 25th

Summary Title : Day #9 (Out of Sim)

Author : Julielynn Wong

Mission Status : All nominal

Sol Activity Summary:

COM continued data collection for the thoracentesis and bike-powered 3D printing studies.

COM + XO continued 3D printing items in Level 1 of the Hab.

ENG + GHO signed out their duties to XO + COM.

ENG + GHO conducted testing of the new drone controller, completed surveys, and submitted qualitative feedback to manufacturer.

XO transported ENG/Astronomer + GHO/Journalist to GJT this PM.

Look ahead plan : Continue 3D printing projects, drone flight mapping, drone controller testing, and data collection for bike-powered 3D printing studies.

Anomalies in work : None

Weather : Sunny

Crew physical status : Fine

EVA : Drone flight testing in PM.

Reports to be filed : Operations, GreenHab, Astronomy

Support Requested : None

Journalist Report – Mar 24th

Journalist Report March 24

Journalist: Diane Rothberg

We are operating as a reduced crew today, as we sadly said goodbye to our two capable, young scientists, Tiffany and Kevin.

We are out of sim, but there is still the Hab to be run, and Tom did his operations duties, I took care of the plants in the Greenhab, and Julielynn compiled scientific data and took care of other administrative loose ends.

Tom and I got to take advantage of this glorious, sunny day by going for a hike. It was a real treat, being back in the “earth’s atmosphere”, and outdoors without the burden of a heavy space suit. The terrain here is awe-inspiring, oddly shaped hills and rock formations in every brown-red hue imaginable, and fossils from thousands of years ago everywhere!

We head out of here tomorrow, but what a privilege it has been to participate in this rare, and wonderful “Mars” experience. Thank you MDRS, and thank you, Medical Makers!

Copyright © The Mars Society. All rights reserved. | Main Site